Gucchi:

Given the function g(x) = −3x + 4, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is smaller than the value of g(4). The value of g(−2) is the same as the value of g(4). The values of g(−2) and g(4) cannot be compared. The value of g(−2) is larger than the value of g(4).

Gucchi:

Am I right with D?

Gucchi:

@axie

Gucchi:

@tetsxpreme

axie:

So for this equation, you insert the given number in exchange for g(x), since you have two equations we need to do this twice

Gucchi:

hmm ok

axie:

g(x) = -3x + 4 g(-2) = -3 x (-2) + 4 = 6 + 4 = 10 g(4) = -3 x 4 + 4 -12 + 4 = = -8

Gucchi:

alr so D it is?

Gucchi:

i understand that

axie:

Yup!

Gucchi:

thank you